# OpenVLThinker: Complex Vision-Language Reasoning via Iterative SFT-RL Cycles

### v2

@@ 1 Introduction
Proprietary large language models (LLMs), notably OpenAI’s o-series [29] and Google’s Gemini-2.5 Pro [17], have demonstrated impressive multi-step reasoning abilities of planning, reflection, and verification. Recent open-weight models [51, 27, 46, 82, 81] (e.g., DeepSeek-R1 [19] and smaller LLMs like S1 [48] and QwQ-32B [61]) show that reinforcement learning (RL) with verifiable rewards effectively reproduces these advanced capabilities, significantly boosting performance on challenging mathematical and logical tasks.
Unlike text-only LLMs, it remains unclear whether open-source large vision-language models (LVLMs) can effectively adopt similar sophisticated reasoning strategies. Modern LVLMs such as LLaVA-NeXT [37] and Qwen2.5-VL [3] benefit from extensive vision-language pretraining and demonstrate strong visual instruction-following capabilities. However, they rarely demonstrate advanced reasoning behaviors like GPT-o1 or DeepSeek-R1.
Moreover, it is known that reasoning capabilities can generally be distilled from larger LLMs to smaller ones through supervised fine-tuning (SFT) on chain-of-thought demonstrations [35, 31] for text-only tasks. This recipe has been recently applied in distills demonstrations from DeepSeek-R1 (LIMO [76], S1 [48] and OpenThinker [60]) followed by optional RL fine-tuning [77]. However, adapting this method to LVLMs does not work. Proprietary LVLMs, such as OpenAI’s o1/o3 and Google’s Gemini, do not expose their internal reasoning paths, making their outputs unsuitable for distillation. Therefore, most recent attempts are focusing on improving LVLMs through distillation from text-only R1 reasoning models (see discussion in Section 2.2). Unfortunately, our experiments show that naively fine-tuning LVLMs on reasoning paths generated from text-based DeepSeek-R1 with image captions leads to a non-trivial performance drop (see Figure 4), primarily due to a lack of precise visual grounding. Similar observations can be found in [6, 74].
@@
    Curriculum RL. In subsequent iterations, we further enhance the LVLM’s reasoning through RL exploration with Group Relative Policy Optimization (GRPO) [54], which splits training into two rounds to form a smooth curriculum.
@@ 2.1 Complex Chain-of-Thought Reasoning
Since the introduction of OpenAI’s O1 model [29], researchers have shown strong interest in reproducing and enhancing the complex reasoning capabilities of LLMs [51, 27, 46, 82, 81], partly due to its superior performance on mathematical benchmarks. [19] introduce the open-source DeepSeek-R1 model and investigate how RL with verifiable rewards can promote advanced chain-of-thought reasoning and reflective behaviors. This development inspired a line of research focused on open-source reproduction [44, 24, 83, 40, 60] and the analysis of such complex reasoning in mathematical problem solving [77, 71, 76, 9]. In parallel, several recent studies have similarly explored the effects of test-time scaling on encouraging more complex model reasoning behaviors [48, 53, 38, 16, 85, 58]. However, the majority of research have significantly advanced text-based reasoning, and development of vision-language reasoning is much more initial.
@@ 2.2 Vision-Language Reasoning Model
Recent advancements in large vision-language models (LVLMs) stem from open-source LLMs [63, 64, 13, 73] and text-aligned image encoders [52, 34]. Integrating these components has enabled LVLMs to follow diverse visual instructions and generate meaningful responses [36, 14, 15, 10, 37, 3]. Parallel to the model development, researchers have also been interested in eliciting CoT reasoning chains from LVLMs via prompting [87, 88, 47, 25] or fine-tuning [20, 72, 62, 12]. These reasoning models remain mostly on a shallow level of common step-by-step prompting, without self-reflections or self-verifications.
Concurrent work. Very recently, many studies have started exploring how to equip LVLMs with R1-like reasoning capabilities through distillation from text-only reasoning models [5, 74, 26, 32] or directly rely on RL [89, 42] for self-exploration. Further advancements [55, 69, 6, 45, 70, 66, 39, 84] have focused on improving performance in visual math reasoning, which marks the transition from early-stage exploration to more effective complex vision-language reasoning. Please note that most of these works are within the two months before the submission date, and some of them do not even have associated technical reports available yet. Our work aligns with these studies and contributes unique insights into the role of SFT for complex reasoning, along with an iterative SFT-RL framework to further advance research in this direction.
@@ 3 Preliminaries
Reinforcement Learning (RL). RL approaches fine-tune LLMs via human preferences modeled under the Bradley-Terry model [49, 11, 54, 1]: $p(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})=\sigma\bigl{(}r(\mathbf{x},\mathbf{y}_{w})-r(\mathbf{x},\mathbf{y}_{l})\bigr{)}$, where $\mathbf{y}_{w}$ and $\mathbf{y}_{l}$ denote preferred and dispreferred responses, respectively, and $\sigma(t)=1/(1+e^{-t})$ is the sigmoid function. The common RL objective under the Bradley-Terry assumption of the reward model $r(\mathbf{x},\mathbf{y})$ is thus
@@
where $\beta>0$ is the KL penalty coefficient. Under this framework, [54] introduced Group Relative Policy Optimization (GRPO) by sampling a group of response trajectories $\{\mathbf{o}_{i}\}_{i=1}^{G}$ from the old policy model $\bm{\theta}_{\mathrm{old}}$ for each query $\mathbf{x}$, with the objective as maximizing:
[…]
@@ 4.1 The Role of SFT and RL
The standard distillation approach used for text-only reasoning cannot be directly applied because the R1 model does not support visual input, and other proprietary LVLMs, such as OpenAI’s o1/o3, do not expose their internal reasoning paths. To learn reasoning behaviors from R1, we instead use the target model as a captioning model, prompting it to generate detailed textual descriptions for each image. Subsequently, these captions serve as proxies for the images when input into a text-based R1 reasoning model, QwQ-32B [61], which then generates $k$ candidate reasoning chains. Among these candidates, we select the shortest reasoning chain that correctly arrives at the final answer to avoid excessive reasoning length after SFT (further details in Section 5.2). The overall procedure is summarized in Figure 5.
@@
Complex reasoning behaviors in LLMs have been described using various terms, including long CoT [77] and aha moments [19]. At their core, these behaviors reflect autonomous planning, reflection, and verification steps that occur during inference. We refer to them as inference-time actions, which are often triggered by specific tokens such as “wait”. To examine how SFT and RL influence these reasoning actions, we identify eight representative keywords corresponding to perception, question comprehension, planning, reflection, and seeking alternatives.
@@
Conversely, direct RL training without prior SFT struggles to efficiently induce reasoning behaviors, exemplified by the absence of reflection keywords (e.g., “wait”) even after an equivalent training volume. Concurrent research by [66], which solely relies on RL, addresses this by explicitly appending relevant keywords during training rollouts. These observations support our argument that SFT plays a critical role in highlighting desirable reasoning actions, providing an efficient and effective foundation for RL to build upon. In contrast, RL primarily serves to further refine and enhance performance.
@@ 4.2 Iterative Improvement
The model obtained after the first iteration (GRPO-Iter1) demonstrates enhanced complex reasoning capabilities and improved reliability in processing visual inputs compared to methods based on image-to-text conversion. This advancement positions GRPO-Iter1 as an effective source for generating higher-quality reasoning demonstrations. Consequently, we propose an iterative self-improvement strategy, inspired by established methodologies such as iterative SFT in ReST-EM [57] and iterative direct preference optimization (DPO) schemes [79, 50], both of which have shown substantial effectiveness in iterative training processes and fall under the Expectation-Maximization framework [57].
Specifically, in each iteration, we sample a new set of enhanced reasoning traces using the model trained in the preceding iteration. These refined demonstrations are then utilized to retrain the base model [^1] To maintain stability, we retrain the model from scratch at each iteration with the newly generated dataset, as similar to some iterative approaches in text-only domain [57, 23]., thereby progressively elevating its reasoning performance. The overall iterative pipeline is illustrated in Figure 2, and the consistent incremental performance gains achieved through successive iterations are depicted in Figure 4.
@@ 4.3 Two-Stage Source-Based Curriculum RL
To ensure effective exploration during reinforcement learning (RL), we assess the difficulty of data sources, aiming to provide data that is challenging yet appropriate for the model’s proficiency level. Specifically, we utilize GPT-4o to rate the difficulty of five representative examples drawn from various data sources such as FigureQA [30], MapQA [4], and GeoQA [7], in a similar fashion to the text-based evaluation in DeepMath-103K [22]. Additionally, we employ the base model, Qwen2.5-VL-7B [^2] In alignment with previous R1 reasoning research [77, 76], we choose the base model from Qwen2.5 family for their strong general capability obtained in pre-training., to obtain its error rates as a complementary difficulty indicator. We standardize independently using z-score normalization for both the GPT-4o rating and base model error rates and compute the average of the two. Based on this composite score, we categorize the data sources into Easy, Medium, and Hard groups via k-means clustering in 1d space. With these categories, we construct two difficulty-specific datasets: $\mathcal{D}_{\mathrm{RL(Medium)}}$ and $\mathcal{D}_{\mathrm{RL(Hard)}}$. Our curriculum training thus proceeds in two stages within one iteration, sequentially training on $\mathcal{D}_{\mathrm{RL(Medium)}}$ and $\mathcal{D}_{\mathrm{RL(Hard)}}$.
@@ 5 Experiments
Training setup. We take Qwen2.5-VL-7B [3] as the base model and perform three iterations of the SFT-RL cycle as illustrated in Section 4, applying full fine-tuning for both SFT and RL. Our training framework is based on LLaMA-Factory [^3] https://github.com/hiyouga/LLaMA-Factory for SFT and EasyR1444https://github.com/hiyouga/EasyR1 for RL. We source our training data from the established LLaVA-OneVision [33] and specifically consider the 14 data sources in overlap with MathV360K [56] (Table 4). Based on our preliminary experiments, we equally draw 500 examples from each source to form the SFT seed dataset of 7K examples, where for each iteration we collect distillation data via rejection sampling, resulting in a final 3K SFT data. We then classify the data sources into easy, medium and hard (as detailed in Table 4). We construct the 3K medium-level RL training data from the 5 sources that we identified as medium difficulty. Finally, we construct 6K hard-level RL training data from the 3 most difficult sources, summing up to 12K data in total for each iteration that trains from the base model. We defer the training hyperparameters to Appendix B.
@@
  • Math reasoning: MathVista [43], MathVerse [86] and MathVision [67]. The three benchmarks evaluate how LVLMs interpret and reason with diagrams in visual math problems through both multiple-choice and free-form questions.
  • General reasoning: MMMU-Pro [80] and EMMA [21]. MMMU-Pro spans 30 subjects across 183 subfields, including business, medicine, and science. EMMA evaluates in physics, chemistry, coding, and math.
  • Perception: HallusionBench [18], designed to evaluate LVLMs’ susceptibility to language hallucination and visual illusion.
Baselines. We evaluate the non-reasoning base model Qwen2.5-VL-7B as a primary baseline to demonstrate the improvements introduced by our method. Additionally, we include the reported performance of proprietary models, including GPT-4o [28] and Claude-3.5-Sonnet [2], alongside open-source LVLMs such as Mulberry-7B [75], InternVL2.5-8B [8], Kimi-VL-16B [59], and Qwen2-VL-7B [68], as reference points. Crucially, to highlight the effectiveness of our iterative SFT-RL training strategy, we compare our model with concurrent approaches employing a single round of SFT distillation and RL at the same model scale (7B), yet utilizing significantly larger training datasets. These concurrent models include R1-VL-7B [84], R1-Onevision-7B [74], and VLAA-Thinker-Qwen2.5VL-7B [6]. Notably, R1-Onevision and VLAA-Thinker-Qwen2.5VL-7B also start from the same base model (Qwen2.5-VL-7B) as ours, using 165K and 150K total data, respectively. In contrast, our model achieves better performance with only 12K training samples from the base model.

### v3

@@ 1 Introduction
Proprietary large language models (LLMs), notably OpenAI’s o-series [jaech2024openai] and Google’s Gemini-2.5 Pro [gemini2.5pro], have demonstrated impressive multi-step reasoning abilities of planning, reflection, and verification. Recent open-weight models [qin2024o1replication, huang2024o1replication, min2024imitateexplore, zhang2024llama, zhang2025rest] (e.g., DeepSeek-R1 [guo2025deepseek] and smaller LLMs like S1 [muennighoff2025s1simple] and QwQ-32B [qwq32b]) show that reinforcement learning (RL) with verifiable rewards effectively reproduces these advanced capabilities, significantly boosting performance on challenging mathematical and logical tasks.
Unlike text-only LLMs, it remains unclear whether open-source large vision-language models (LVLMs) can effectively adopt similar sophisticated reasoning strategies. Modern LVLMs such as LLaVA-NeXT [liu2024llavanext] and Qwen2.5-VL [bai2025qwen2] benefit from extensive vision-language pretraining and demonstrate strong visual instruction-following capabilities. However, they rarely demonstrate advanced reasoning behaviors like GPT-o1 or DeepSeek-R1.
Moreover, it is known that reasoning capabilities can generally be distilled from larger LLMs to smaller ones through supervised fine-tuning (SFT) on chain-of-thought demonstrations [li2023symbolic, kim-etal-2023-cot] for text-only tasks. This recipe has been recently applied in distills demonstrations from DeepSeek-R1 (LIMO [ye2025limoreasoning], S1 [muennighoff2025s1simple] and OpenThinker [openthoughts]) followed by optional RL fine-tuning [yeo2025demystifying]. However, adapting this method to LVLMs does not work. Proprietary LVLMs, such as OpenAI’s o1/o3 and Google’s Gemini, do not expose their internal reasoning paths, making their outputs unsuitable for distillation. Therefore, most recent attempts are focusing on improving LVLMs through distillation from text-only R1 reasoning models (see discussion in Section 2.2). Unfortunately, our experiments show that naively fine-tuning LVLMs on reasoning paths generated from text-based DeepSeek-R1 with image captions leads to a non-trivial performance drop (see Figure 4), primarily due to a lack of precise visual grounding. Similar observations can be found in [chen2025sft, yang2025r1onevision].
@@
    Curriculum RL. In subsequent iterations, we further enhance the LVLM’s reasoning through RL exploration with Group Relative Policy Optimization (GRPO) [shao2024deepseekmath], which splits training into two rounds to form a smooth curriculum.
@@ 2.1 Complex Chain-of-Thought Reasoning
Since the introduction of OpenAI’s O1 model [jaech2024openai], researchers have shown strong interest in reproducing and enhancing the complex reasoning capabilities of LLMs [qin2024o1replication, huang2024o1replication, min2024imitateexplore, zhang2024llama, zhang2025rest], partly due to its superior performance on mathematical benchmarks. [guo2025deepseek] introduce the open-source DeepSeek-R1 model and investigate how RL with verifiable rewards can promote advanced chain-of-thought reasoning and reflective behaviors. This development inspired a line of research focused on open-source reproduction [deepscaler2025, OpenReasonerZero2025, zhang2025dpor1, liu2025oatzero, openthoughts] and the analysis of such complex reasoning in mathematical problem solving [yeo2025demystifying, xie2025logicrl, ye2025limoreasoning, chen2025empirical]. In parallel, several recent studies have similarly explored the effects of test-time scaling on encouraging more complex model reasoning behaviors [muennighoff2025s1simple, setlur2025scaling, liu20251bllmsurpass405b, geiping2025scaling, zhang2024generative, singhi2025solve]. However, the majority of research have significantly advanced text-based reasoning, and development of vision-language reasoning is much more initial.
@@ 2.2 Vision-Language Reasoning Model
Recent advancements in large vision-language models (LVLMs) stem from open-source LLMs [touvron2023llama, touvron2023llama2, dubey2024llama, yang2024qwen2] and text-aligned image encoders [radford2021learning, li2022grounded]. Integrating these components has enabled LVLMs to follow diverse visual instructions and generate meaningful responses [liu2023visual, gao2023llamaadapterv2, gao2024sphinx, dai2024instructblip, liu2024llavanext, bai2025qwen2]. Parallel to the model development, researchers have also been interested in eliciting CoT reasoning chains from LVLMs via prompting [zhang2023multimodal, zheng2023ddcot, mondal2024kam, hu2024visual] or fine-tuning [guo2024mammoth, xu2024llava, thawakar2025llamav, du2025virgo]. These reasoning models remain mostly on a shallow level of common step-by-step prompting, without self-reflections or self-verifications.
Concurrent work. Very recently, many studies have started exploring how to equip LVLMs with R1-like reasoning capabilities through distillation from text-only reasoning models [vl-thinking2025, yang2025r1onevision, huang2025visionr1, MMR1-Math2025] or directly rely on RL [zhou2025r1, liu2025visual] for self-exploration. Further advancements [shen2025vlm, wang2025visualprm, chen2025sft, meng2025mm, wang2025sota, wang2025vl, liu2025noisyrollout, zhang2025r1] have focused on improving performance in visual math reasoning, which marks the transition from early-stage exploration to more effective complex vision-language reasoning. Please note that most of these works are within the two months before the submission date, and some of them do not even have associated technical reports available yet. Our work aligns with these studies and contributes unique insights into the role of SFT for complex reasoning, along with an iterative SFT-RL framework to further advance research in this direction.
@@ 3 Preliminaries
Reinforcement Learning (RL). RL approaches fine-tune LLMs via human preferences modeled under the Bradley-Terry model [ouyang2022training, dong2024rlhf, shao2024deepseekmath, ahmadian2024back]: $p(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})=\sigma\bigl(r(\mathbf{x},\mathbf{y}_{w})-r(\mathbf{x},\mathbf{y}_{l})\bigr)$, where $\mathbf{y}_{w}$ and $\mathbf{y}_{l}$ denote preferred and dispreferred responses, respectively, and $\sigma(t)=1/(1+e^{-t})$ is the sigmoid function. The common RL objective under the Bradley-Terry assumption of the reward model $r(\mathbf{x},\mathbf{y})$ is thus
@@
where $\beta>0$ is the KL penalty coefficient. Under this framework, [shao2024deepseekmath] introduced Group Relative Policy Optimization (GRPO) by sampling a group of response trajectories $\{\mathbf{o}_{i}\}_{i=1}^{G}$ from the old policy model $\bm{\theta}_{\mathrm{old}}$ for each query $\mathbf{x}$, with the objective as maximizing:
$\mathbb{E}\Bigl[\!\frac{1}{G}\!\sum_{i=1}^{G}\,\frac{1}{\lvert\mathbf{o}_{i}\rvert}\sum_{t=1}^{\lvert\mathbf{o}_{i}\rvert}\min\Bigl(\,\frac{p_{\bm{\theta}}(o_{i,t}\mid\mathbf{x},\,\mathbf{o}_{i,<t})}{p_{\bm{\theta}_{\mathrm{old}}}(o_{i,t}\mid\mathbf{x},\,\mathbf{o}_{i,<t})}\,\widehat{A}_{i,t},\;\!\!\mathrm{clip}\Bigl(\frac{p_{\bm{\theta}}(o_{i,t}\mid\mathbf{x},\,\mathbf{o}_{i,<t})}{p_{\bm{\theta}_{\mathrm{old}}}(o_{i,t}\mid\mathbf{x},\,\mathbf{o}_{i,<t})},\,1-\epsilon,\,1+\epsilon\Bigr)\,\widehat{A}_{i,t}\Bigr)\Bigr]$
@@ 4.1 The Role of SFT and RL
The standard distillation approach used for text-only reasoning cannot be directly applied because the R1 model does not support visual input, and other proprietary LVLMs, such as OpenAI’s o1/o3, do not expose their internal reasoning paths. To learn reasoning behaviors from R1, we instead use the target model as a captioning model, prompting it to generate detailed textual descriptions for each image. Subsequently, these captions serve as proxies for the images when input into a text-based R1 reasoning model, QwQ-32B [qwq32b], which then generates $k$ candidate reasoning chains. Among these candidates, we select the shortest reasoning chain that correctly arrives at the final answer to avoid excessive reasoning length after SFT (further details in Section 5.2). The overall procedure is summarized in Figure 5.
@@
Complex reasoning behaviors in LLMs have been described using various terms, including long CoT [yeo2025demystifying] and aha moments [guo2025deepseek]. At their core, these behaviors reflect autonomous planning, reflection, and verification steps that occur during inference. We refer to them as inference-time actions, which are often triggered by specific tokens such as “wait”. To examine how SFT and RL influence these reasoning actions, we identify eight representative keywords corresponding to perception, question comprehension, planning, reflection, and seeking alternatives.
@@
Conversely, direct RL training without prior SFT struggles to efficiently induce reasoning behaviors, exemplified by the absence of reflection keywords (e.g., “wait”) even after an equivalent training volume. Concurrent research by [wang2025vl], which solely relies on RL, addresses this by explicitly appending relevant keywords during training rollouts. These observations support our argument that SFT plays a critical role in highlighting desirable reasoning actions, providing an efficient and effective foundation for RL to build upon. In contrast, RL primarily serves to further refine and enhance performance.
@@ 4.2 Iterative Improvement
The model obtained after the first iteration (GRPO-Iter1) demonstrates enhanced complex reasoning capabilities and improved reliability in processing visual inputs compared to methods based on image-to-text conversion. This advancement positions GRPO-Iter1 as an effective source for generating higher-quality reasoning demonstrations. Consequently, we propose an iterative self-improvement strategy, inspired by established methodologies such as iterative SFT in ReST-EM [singh2023beyond] and iterative direct preference optimization (DPO) schemes [yuan2024selfrewardinglanguagemodels, pang2024iterative], both of which have shown substantial effectiveness in iterative training processes and fall under the Expectation-Maximization framework [singh2023beyond].
Specifically, in each iteration, we sample a new set of enhanced reasoning traces using the model trained in the preceding iteration. These refined demonstrations are then utilized to retrain the base model [^1] To maintain stability, we retrain the model from scratch at each iteration with the newly generated dataset, as similar to some iterative approaches in text-only domain [singh2023beyond, hosseini2024v]., thereby progressively elevating its reasoning performance. The overall iterative pipeline is illustrated in Figure 2, and the consistent incremental performance gains achieved through successive iterations are depicted in Figure 4.
@@ 4.3 Two-Stage Source-Based Curriculum RL
To ensure effective exploration during reinforcement learning (RL), we assess the difficulty of data sources, aiming to provide data that is challenging yet appropriate for the model’s proficiency level. Specifically, we utilize GPT-4o to rate the difficulty of five representative examples drawn from various data sources such as FigureQA [kahou2017figureqa], MapQA [chang2022mapqa], and GeoQA [chen2021geoqa], in a similar fashion to the text-based evaluation in DeepMath-103K [he2025deepmath]. Additionally, we employ the base model, Qwen2.5-VL-7B [^2] In alignment with previous R1 reasoning research [yeo2025demystifying, ye2025limoreasoning], we choose the base model from Qwen2.5 family for their strong general capability obtained in pre-training., to obtain its error rates as a complementary difficulty indicator. We standardize independently using z-score normalization for both the GPT-4o rating and base model error rates and compute the average of the two. Based on this composite score, we categorize the data sources into Easy, Medium, and Hard groups via k-means clustering in 1d space. With these categories, we construct two difficulty-specific datasets: $\mathcal{D}_{\mathrm{RL(Medium)}}$ and $\mathcal{D}_{\mathrm{RL(Hard)}}$. Our curriculum training thus proceeds in two stages within one iteration, sequentially training on $\mathcal{D}_{\mathrm{RL(Medium)}}$ and $\mathcal{D}_{\mathrm{RL(Hard)}}$.
@@ 5 Experiments
Training setup. We take Qwen2.5-VL-7B [bai2025qwen2] as the base model and perform three iterations of the SFT-RL cycle as illustrated in Section 4, applying full fine-tuning for both SFT and RL. Our training framework is based on LLaMA-Factory [^3] https://github.com/hiyouga/LLaMA-Factory for SFT and EasyR1444https://github.com/hiyouga/EasyR1 for RL. We source our training data from the established LLaVA-OneVision [li2024llava] and specifically consider the 14 data sources in overlap with MathV360K [shi2024math] (Table 4). Based on our preliminary experiments, we equally draw 500 examples from each source to form the SFT seed dataset of 7K examples, where for each iteration we collect distillation data via rejection sampling, resulting in a final 3K SFT data. We then classify the data sources into easy, medium and hard (as detailed in Table 4). We construct the 3K medium-level RL training data from the 5 sources that we identified as medium difficulty. Finally, we construct 6K hard-level RL training data from the 3 most difficult sources, summing up to 12K data in total for each iteration that trains from the base model. We defer the training hyperparameters to Appendix C.
@@
  • Math reasoning: MathVista [lu2023mathvista], MathVerse [zhang2024mathverse] and MathVision [wang2024measuring]. The three benchmarks evaluate how LVLMs interpret and reason with diagrams in visual math problems through both multiple-choice and free-form questions.
  • General reasoning: MMMU-Pro [yue2024mmmu] and EMMA [hao2025can]. MMMU-Pro spans 30 subjects across 183 subfields, including business, medicine, and science. EMMA evaluates in physics, chemistry, coding, and math.
  • Perception: HallusionBench [guan2024hallusionbench], designed to evaluate LVLMs’ susceptibility to language hallucination and visual illusion.
Baselines. We evaluate the non-reasoning base model Qwen2.5-VL-7B as a primary baseline to demonstrate the improvements introduced by our method. Additionally, we include the reported performance of proprietary models, including GPT-4o [hurst2024gpt] and Claude-3.5-Sonnet [claude], alongside open-source LVLMs such as Mulberry-7B [yao2024mulberry], InternVL2.5-8B [chen2024expanding], Kimi-VL-16B [team2025kimi], and Qwen2-VL-7B [wang2024qwen2], as reference points. Crucially, to highlight the effectiveness of our iterative SFT-RL training strategy, we compare our model with concurrent approaches employing a single round of SFT distillation and RL at the same model scale (7B), yet utilizing significantly larger training datasets. These concurrent models include R1-VL-7B [zhang2025r1], R1-Onevision-7B [yang2025r1onevision], and VLAA-Thinker-Qwen2.5VL-7B [chen2025sft]. Notably, R1-Onevision and VLAA-Thinker-Qwen2.5VL-7B also start from the same base model (Qwen2.5-VL-7B) as ours, using 165K and 150K total data, respectively. In contrast, our model achieves better performance with only 12K training samples from the base model.